□ Map to the Conference Site (Brest State Technical Univ.)

• If you are lost, try Russian! ⋆ “Gje (= Where) Politeknicheski Institut?” ⋯ on the street ⋆ “Na (= To) Politeknicheski Institut?” ⋯ in taxi.